# CoRaiS: Lightweight Real-Time Scheduler for Multi-Edge Cooperative Computing

### v1

@@ IV-B Training CoRaiS
We generate synthetic data to build the training dataset $D$ for some reasons. Firstly, even though Alibaba has provided the most prominent cluster dataset, CoRaiS cannot be trained using the dataset, because majority of services in the dataset are related with online shopping applications, which are not ideal services [^3] We explain ideal services in Section III-C. Ideal services refers to services that has a functional relationship between running time and data size. focused on in this paper. Only a few offline services may meet the characteristic of ideal services, but it is difficult to separate them from the dataset. Moreover, the dataset is closely tied to the specific multi-edge cooperative system built by Alibaba, hence, there is uncertainty regarding whether a scheduling policy trained on the dataset would perform well on other multi-edge systems. Secondly, no other datasets containing enough data can be used to train CoRaiS. Thirdly, when generating synthetic data, we can vary the multi-edge cooperative environments and the scheduled services. In this way, the synthetic dataset becomes more comprehensive and better supports CoRaiS in learning a more versatile scheduling policy. To introduce variation of multi-edge cooperative environments, we can vary the number and position of heterogeneous edges, as well as the number of requests submitted to the each edges. To diversify the scheduled services, we can establish different computation time estimation function and service replica numbers (explained in Section III-C1) to express runtime performance of different services.
IV-C Decode Strategies of CoRaiS
@@ V Evaluation
To demonstrate that CoRaiS is able to learn a strong policy and give a real-time decision for the multi-edge scheduling problem, we designs three experiments: conventional test, generalization test, and characteristic validation. Conventional test refers to evaluating the performance of CoRaiS on a dataset that matches the same scale [^4] In this paper, the scale of dataset refers to the size of the multi-edge cooperative system, including the number of edges in the system and the number of requests submitted to it. as the one it was trained on. Generalization test refers to evaluating the performance of CoRaiS on a dataset that has large scale than the one it was originally trained on. Characteristic validation is used to assess whether CoRaiS can perceive the workload and heterogeneity of edges, and autonomously implement load balancing.
TABLE II: Conventional Test Results
(Time(s) and Gap: the lower the better. EN and RN denote the number of edges and new requests respectively. 1k=1000.
$(x\%)$ in the column of Time indicates $x\%$ of test instances are resolved. $(y)$ in the column Gap indicates that only successfully resolved instances are counted while computing quality difference.)
EN=5, RN=50 EN=10, RN=50 EN=5, RN=100 EN=10, RN=100
Method Time(s) Gap Time(s) Gap Time(s) Gap Time(s) Gap
Gurobi(10s) 0.3592 1 0.3473 1 0.6542 1 3.4865 1
Gurobi(1s) 0.3252 1 0.3373 1 0.1252 1.0001 0.4921 1.0001
Local - 2.1712 - 2.8032 - 1.8594 - 2.7055
Random(1) - 2.1987 - 2.8023 - 1.9198 - 2.759
Random(100) - 1.2916 - 1.8148 - 1.2732 - 1.7422
Random(1k) - 1.2916 - 1.6282 - 1.1863 - 1.5854
FC1-CoRaiS(greedy) 0.0050 3.1129 0.0052 4.638 0.0052 3.346 0.0052 5.3293
FC2-CoRaiS(greedy) 0.0051 4.0758 0.0051 5.7307 0.0052 4.4914 0.0051 7.8728
FC3-CoRaiS(greedy) 0.0051 4.0758 0.0052 5.7307 0.0053 4.4914 0.0051 7.8728
CoRaiS(greedy) 0.0051 1.0783 0.0052 1.0953 0.0052 1.0450 0.0051 1.1083
FC1-CoRaiS(100) 0.0051 1.4118 0.0051 1.8141 0.0052 1.2704 0.0051 1.7481
FC2-CoRaiS(100) 0.0051 1.4131 0.0052 1.8163 0.0051 1.2703 0.0052 1.7491
FC3-CoRaiS(100) 0.0051 1.4124 0.0052 1.8182 0.0052 1.2704 0.0051 1.7463
CoRaiS(100) 0.0051 1.0221 0.0052 1.0267 0.0052 1.0091 0.0051 1.0361
FC1-CoRaiS(1k) 0.0050 1.2956 0.0052 1.6274 0.0052 1.1837 0.0053 0.8769
FC2-CoRaiS(1k) 0.0051 1.2913 0.0052 1.6269 0.0052 1.1857 0.0051 1.5858
FC3-CoRaiS(1k) 0.0051 1.2929 0.0051 1.6276 0.0052 1.1856 0.0051 1.5851
CoRaiS(1k) 0.0051 1.0148 0.0052 1.0186 0.0052 1.0069 0.0051 1.0297
Gurobi(0.005s) (44.81%) (1.7981) (19.69%) (2.6884) (6.02%) (1.184) (5.46%) -
Gurobi(0.01s) (83.82%) (1.2584) (45.91%) (2.0371) (34.91%) (1.3904) (3.40%) -
Gurobi(0.05s) 0.0584 1.0009 0.0886 1.0026 0.0629 1.0013 0.1211 1.0334
V-A Experiment Descriptions
@@
Instance generation   The training and testing datasets are generated as the same rules. Given the number of edges and requests $(Q,Z)$, for each edge $e_{q}$, the coordinates are randomly sampled under the uniform distribution in $(0,1)^{2}$; the supported maximum service replica number $rn_{q}$ is randomly sampled in $\{1,2,3,4\}$. The functional relationship between computing time and the size of input data packets is modeled as linear functions [^5] Without loss of generality, we model the relationship using the linear functions. But in practical, other relationships, such as quadratic function, are allowed to train CoRaiS. for all edges during simulations. Moreover, to represent the heterogeneity across edges, different coefficients are randomly sampled from a uniform distribution within the range $(0,1)$.
@@
To make CoRaiS can learn to adopt to any initial system-level state, we randomly generate some requests as backlogs for each edge while generating training dataset and testing dataset. The number of backlogs in $Q_{q}^{le}$ and $Q_{q}^{in}$ of edge $e_{q}$ are randomly sampled from $(0,100)$. For any backlog $r_{x}$ in $Q_{q}^{le}$, its input data size $f_{x}$ is uniformly sampled from $(0,1)$. For any backlog $r_{k}$ in $Q_{q}^{in}$, its source edge $\varpi_{k}$ is sampled from $[Q]-\{q\}$, and its input data size $f_{k}$ is uniformly sampled from $(0,1)$. With these backlogs, the service-oriented workload evaluation are carried out as (1)-(3).
@@
For any new request $r_{z}$ that will be scheduled in current period, its source edge $\varpi_{z}$ is sampled from $[Q]$, and the size of related input data $f_{z}$ is uniformly sampled from $(0,1)$.
@@
Hyperparameters   Learnable parameters are initialized as Uniform($-1/\sqrt{d},1/\sqrt{d}$), with $d$ is the input dimension; learning rate $lr=1e-5$; batch size is 128, $S=64$ while using $S$-samples batch RL; $C_{1}=10$ and $C_{2}=0.5$. The edge encoder and request encoder have $L=5$ and $K=3$ attention layers, respectively. $\mathbf{MHA}$ and $\mathbf{FC}$ in two embedding modules same structure, i.e. $\mathbf{MHA}$ has 8 heads, $\mathbf{FC}$ has one hidden sublayer with dimension 512 and ReLu activation.
@@
Baselines   It must be mentioned that as our best knowledge, CoRaiS is the first artificial intelligence based work that breaks many assumptions of prior works, deals with a multi-edge computing system where every edge is allowed to receive requests from clients, and optimizes an objective function that minimize the responding time over all requests. In order to evaluate the performance of CoRaiS, the results generated by CoRaiS are compared with that produced by exact solver (Gurobi). Gurobi is able to help calculate exact solutions. Meanwhile, we propose two heuristic approaches (Local and Random) as the heuristic baselines in the scenario. Three learning-based baselines are designed as well.
@@
  • Heuristic baselines: two algorithms are used to provide heuristic baselines. (i) Local: executing all requests at their source locations. (ii) Random: randomly sampling the execution edges for all requests. It is allowed to sample multiple times and report the best one.
@@
  • Learning-based baselines (ablation studies for aligning modules of CoRais): three synthetic neural network models are used to illustrate the effectiveness of CoRaiS. (i) FC1-CoRaiS: replacing multi-head attention aligning mechanism of edge encoder in CoRaiS with multi-layer perceptron, and maintaining the same number of neuron parameters. (ii) FC2-CoRaiS: adopting the similar MoD architecture with CoRaiS, but using multi-layer perceptron to replace the multi-head attention aligning mechanism in request encoder. (iii) FC3-CoRaiS: adopting MoD structure, but without aligning mechanisms in both edge encoder and request encoder, only using multi-layer perceptron to embed edges features and requests features. Since FC1-CoRaiS, FC2-CoRaiS and FC3-CoRaiS adopt the similar MoD architecture and have the same input/output with CoRaiS, they are able to use the same decoding strategies with CoRaiS as well.
TABLE III: Generalization Test Results
@@
(CoRaiS is trained on small-scale instances (EN=10, RN=100) and directly applied into larger-scale instances. Time(s) and Gap: the lower the better).
EN=10, RN=200 EN=30, RN=400 EN=50, RN=600 EN=50, RN=800
Method Time(s) Gap Time(s) Gap Time(s) Gap Time(s) Gap
Gurobi(10s) 2.8447 1 3.6902 1 4.0671 1 7.2515 1
Gurobi(0.01s) (0.09%) - 0% 0% 0%
Gurobi(0.02s) (2.99%) - 0% 0% 0%
Gurobi(0.2s) 0.3235 1.0021 (9.89%) - 0% 0%
CoRaiS(greedy) 0.0077 1.3107 0.0119 1.6025 0.0162 1.8897 0.0165 2.3572
CoRaiS(1k) 0.0074 1.0867 0.0118 1.09 0.0159 1.0793 0.0161 1.1954
CoRaiS(10k) 0.0074 1.0752 0.0119 1.0788 0.0161 1.0683 0.0166 1.1720
TABLE IV: Characteristics Validation Results
@@
(CoRaiS is trained on scheduling problems (EN=5 and RN=100).    CoRaiS(1k) is used to sample near-optimal solutions. LB: Load Balancing; WP: Workload Perception; HA: Heterogeneity Awareness; EReqN: the number of requests that are executed at each edges; LCost: the corresponding response time.
@@
A 22.2094 2.7996 11.4865 2.9964 13.1378 2.0203
B 19.4471 2.8274 18.8379 2.9035 14.9200 2.0285
C 19.4511 2.8267 21.5947 2.7351 19.4577 1.9963
D 19.4523 2.8274 23.2921 2.5227 25.7648 1.8249
E 19.4399 2.8259 24.7888 2.1630 26.7197 1.2267
Performance indexes   Two indexes are used to evaluate performance. (i) Time(s): time taken to make scheduling decisions, because the multi-edge cooperative computing system needs an approach that can provide a real-time scheduling decision. (ii) Gap: quality difference of solutions that are generated by CoRaiS and other baselines, compared with the solution generated by Gurobi(10s) in the same instance. Solutions from Gurobi(10s) are considered as the optimal benchmark in the simulations. The gap is computed by (22). $\aleph$ includes CoRaiS and other baselines. $\pi$ signifies the best solution obtained from the specific approach $b$ ($b\in\aleph$). $\hat{\pi}$ refers to the best solution generated by $Gurobi(10s)$. $L(\pi)$ refers to the predicted response time for all requests in the multi-edge cooperative computing system, and $L(\pi)$ is computed by (19). The gap is critical because the multi-edge cooperative computing system also require the real-time approach to produce a high-quality solution.
V-B Results Analysis
V-B1 Analysis of Conventional Test Results

### v2

@@ VI-B Training CoRaiS
VI-C Decoding Strategies of CoRaiS
@@ VII Simulation Evaluation
To demonstrate that CoRaiS is able to learn a strong policy and give a real-time decision for the multi-edge scheduling problem, we designs three experiments: conventional test, generalization test, and characteristic validation. The conventional test refers to evaluating the performance of CoRaiS on a dataset that matches the same scale [^3] In this paper, the scale of dataset refers to the size of the multi-edge cooperative system, including the number of edges in the system and the number of requests submitted to it. as the one it was trained on. The generalization test refers to evaluating the performance of CoRaiS on a dataset that has large scales than the one it was originally trained on. The characteristic validation is used to assess whether CoRaiS can perceive the workload and heterogeneity of edges, and autonomously implement load balancing.
TABLE III: Normal Test Results
CoRaiS is trained and tested on the same small-scale instances $\{(10,5,50),(10,5,100),(20,10,50),(20,10,100)\}$.
Gap-M, Time-M and Time-S: the lower the better.
(10, 5, 50) (10, 5, 100) (20, 10, 50) (20, 10, 100)
Methods Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S
Gurobi(10s) 1.0000 0.4294 1.1454 1.0000 1.3126 1.9815 1.0000 0.1970 0.0170 1.0000 0.5791 1.3249
Local 1.9602 0.0049 0.0008 2.7110 0.0047 0.0006 1.5214 0.0085 0.0008 2.0253 0.0081 0.0015
Predicted 1.9752 0.0090 0.0015 2.7182 0.0123 0.0013 1.5268 0.0125 0.0011 2.0256 0.0155 0.0028
Random(1) 2.0540 0.0049 0.0008 2.7707 0.0048 0.0005 1.5706 0.0086 0.0008 2.0943 0.0082 0.0015
 Random(100) 1.4199 0.0053 0.0008 1.9525 0.0053 0.0006 1.1211 0.0091 0.0008 1.4544 0.0089 0.0016
Random(1k) 1.2962 0.0066 0.0011 1.7691 0.0079 0.0011 1.0624 0.0104 0.0008 1.3342 0.0112 0.0020
FC1-CoRaiS(greedy) 1.4953 0.0135 0.0019 1.5857 0.0132 0.0016 1.4443 0.0171 0.0015 1.5014 0.0163 0.0029
FC2-CoRaiS(greedy) 1.1738 0.0151 0.0034 1.3871 0.0145 0.0015 1.0504 0.0175 0.0025 1.1479 0.0175 0.0031
FC3-CoRaiS(greedy) 1.1802 0.0118 0.0017 1.3422 0.0115 0.0012 1.0345 0.0150 0.0018 1.1279 0.0147 0.0025
CoRaiS(greedy) 1.0476 0.0159 0.0022 1.0823 0.0155 0.0015 1.0182 0.0193 0.0016 1.0409 0.0187 0.0031
FC1-CoRaiS(100) 1.2465 0.0138 0.0019 1.3366 0.0134 0.0016 1.2552 0.0170 0.0015 1.2291 0.0165 0.0029
FC2-CoRaiS(100) 1.0544 0.0150 0.0034 1.2422 0.0146 0.0015 1.0072 0.0181 0.0025 1.0266 0.0179 0.0031
FC3-CoRaiS(100) 1.0566 0.0120 0.0017 1.1949 0.0117 0.0012 1.0038 0.0152 0.0018 1.0259 0.0150 0.0025
CoRaiS(100) 1.0160 0.0157 0.0023 1.0353 0.0154 0.0016 1.0000 0.0192 0.0016 1.0051 0.0185 0.0031
FC1-CoRaiS(1k) 1.1957 0.0138 0.0019 1.2757 0.0135 0.0016 1.2229 0.0166 0.0015 1.1805 0.0166 0.0029
FC2-CoRaiS(1k) 1.0426 0.0151 0.0034 1.2167 0.0146 0.0015 1.0030 0.0183 0.0025 1.0193 0.0179 0.0031
FC3-CoRaiS(1k) 1.0446 0.0122 0.0017 1.1750 0.0117 0.0012 1.0019 0.0156 0.0018 1.0178 0.0151 0.0025
CoRaiS(1k) 1.0113 0.0159 0.0022 1.0274 0.0153 0.0018 1.0000 0.0192 0.0018 1.0036 0.0185 0.0032
Gurobi(0.02s) (0.8%) 0.0772 0.0033 (0%) - - (0%) - - (0%) - -
Gurobi(0.05s) (78.61%) 0.1064 0.0093 (3.2%) 0.1760 0.0162 (2.81%) 0.1713 0.0092 (0%) - -
Gurobi(1s) 1.0001 0.2489 0.3041 1.0015 0.5938 0.4399 1.0000 0.2115 0.0176 1.0005 0.4231 0.1432
Gurobi(5s) 1.0000 0.3963 0.8883 1.0000 1.1848 1.4795 1.0000 0.1991 0.0172 1.0000 0.4875 0.6733
VII-A Experiment Descriptions
@@
Instance generation   The training and testing datasets are generated as the same rules. Given the number of edges in the whole multi-edge cooperative computing system, denoted as $N$, the number of edges in an individual control region $Q(Q<N)$ and the number of requests $Z$, for each edge $e_{q}(q\in\{1,...,N\})$, the coordinates are randomly sampled under the uniform distribution in $(0,1)^{2}$; the supported maximum service replica number $rn_{q}$ is randomly sampled in $\{1,2,3,4\}$. $C_{t}=3$. The functional relationship between computing time and the size of input data packets is modeled as linear functions [^4] Without loss of generality, we model the relationship using the linear functions. But in practical, other relationships, such as quadratic function, are allowed to train CoRaiS. for all edges during simulations. Moreover, to represent the heterogeneity across edges, different coefficients are randomly sampled from a uniform distribution within the range $(0,1)$.
@@
To make CoRaiS can learn to adopt to any initial system-level states, we randomly generate some requests as backlogs for each edge while generating training datasets and testing datasets. The numbers of backlogs in $Q_{q}^{le}$ and $Q_{q}^{in}$ of edge $e_{q}$ are randomly sampled from $(0,100)$. For any backlog $r_{x}$ in $Q_{q}^{le}$, its input data size $f_{x}$ is uniformly sampled from $(0,1)$, output data size $u_{x}$ is uniformly sampled from $(0,0.1)$, and its predicted edge $\delta_{x}$ is sampled from $[N]$. For any backlog $r_{k}$ in $Q_{q}^{in}$, its source edge $\varpi_{k}$ is sampled from $[Q]-\{q\}$, its predicted edge $\delta_{k}$ is sampled from $[N]$, and its input and output data sizes are uniformly sampled from $(0,1)$ and $(0,0.1)$. With these backlogs, the service-oriented workload evaluation is carried out as (1)-(3).
@@
For any new request $r_{z}$ that will be scheduled in current period, its source edge $\varpi_{z}$ is sampled from $[Q]$, its predicted edge $\delta_{z}$ is sampled from $[N]$, and the related input data size $f_{z}$ and output data size $u_{z}$ are uniformly sampled from $(0,1)$ and $(0,0.1)$, respectively.
@@
Hyperparameters   Learnable parameters are initialized as Uniform($-1/\sqrt{d},1/\sqrt{d}$), with $d$ is the input dimension; learning rate $lr=1e-5$; batch size is 128, $S=64$ while using $S$-samples batch RL; $C_{1}=10$ and $C_{2}=0.5$. The edge encoder and request encoder have $L=5$ and $K=3$ attention layers, respectively. $\mathbf{MHA}$ and $\mathbf{FC}$ in two embedding modules have same structure. $\mathbf{MHA}$ has 8 heads, $\mathbf{FC}$ has one hidden sublayer with dimension 512 and ReLu activation.
@@
Baselines   It must be mentioned that as our best knowledge, CoRaiS is the first artificial intelligence based work that breaks many assumptions of prior works, deals with a multi-edge computing system where every edge is allowed to receive requests from clients, and optimizes an objective function that minimize the responding time over all requests. In order to evaluate the performance of CoRaiS, the results generated by CoRaiS are compared with that produced by the exact solver Gurobi. Gurobi is able to calculate exact solutions. Meanwhile, we propose three heuristic approaches (Local, Random and Predicted) as the lightweight scheduling baselines in the scenarios. Three learning-based baselines are designed as well.
@@
  • Lightweight heuristic baselines: two algorithms are used to provide heuristic baselines. (i) Local: executing all requests at their source locations. (ii) Random: randomly sampling the execution edges for all requests. It is allowed to sample multiple times and report the best one. (iii) Predicted: executing all requests at their predicted edges, but if the predicted edges of some requests are out of the control region, the requests will be executed at their source edges.
@@
  • Learning-based baselines (ablation studies for aligning modules of CoRais): three synthetic neural network models are used to illustrate the effectiveness of CoRaiS. (i) FC1-CoRaiS: replacing the multi-head attention aligning mechanism of edge encoder in CoRaiS with the multi-layer perceptron, and maintaining the same number of neuron parameters. (ii) FC2-CoRaiS: adopting the similar MoD architecture with CoRaiS, but using the multi-layer perceptron to replace the multi-head attention aligning mechanism in request encoder. (iii) FC3-CoRaiS: adopting MoD structure, but without aligning mechanisms in both edge encoder and request encoder, only using multi-layer perceptron to embed edges’ features and requests’ features. Since FC1-CoRaiS, FC2-CoRaiS and FC3-CoRaiS adopt the similar MoD architecture and have the same input/output with CoRaiS, they are able to use the same decoding strategies with CoRaiS as well.
TABLE IV: Generalization Test Results on Medium-scale Instances.
@@
CoRaiS is trained on small-scale instances (10, 5, 100) and directly applied into medium-scale instances. $(x\%)$ refers to $x\%$ of the instances are solved.
Gap-M, Time-M and Time-S: the lower the better.
(20, 10, 200) (60, 30, 400) (100, 50, 600) (100, 50, 800)
Methods Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S Gap-M Time-M Time-S
Gurobi(10s) 1.0000 2.0173 3.0761 1.0000 6.1590 1.1698 (97.60%) 16.7801 1.2889 (58.00%) 20.0647 0.7842
Local 3.0285 0.0088 0.0008 2.4070 0.0243 0.0019 2.2841 0.0394 0.0028 2.4602 0.0394 0.0024
Predicted 3.0860 0.0265 0.0018 2.4735 0.0620 0.0032 2.3190 0.0968 0.0047 2.4820 0.1145 0.0042
Random(1k) 1.9392 0.0147 0.0013 1.6445 0.0362 0.0024 1.5753 0.0574 0.0034 1.7049 0.0608 0.0034
CoRaiS(greedy) 1.0739 0.0206 0.0014 1.0139 0.0373 0.0020 0.9822 0.0533 0.0030 0.9400 0.0523 0.0025
CoRaiS(100) 1.0246 0.0201 0.0016 1.0002 0.0362 0.0021 0.9672 0.0523 0.0033 0.8999 0.0520 0.0028
CoRaiS(1k) 1.0202 0.0202 0.0015 1.0002 0.0357 0.0022 0.9672 0.0519 0.0032 0.8999 0.0514 0.0028
Gurobi(0.05s) (0%) - - (0%) - - (0%) - - (0%) - -
Gurobi(1s) 1.0062 0.9511 0.2605 (0%) - - (0%) - - (0%) - -
Gurobi(5s) 1.0006 1.4622 1.5466 1.0032 6.0602 0.8031 (1.2%) 12.4700 0.4302 (0%) - -
Performance indexes   Three indexes are used to evaluate performance.
  • Gap-M: average quality difference of solutions that are generated by CoRaiS and other baselines, compared with the solution generated by Gurobi(10s) in the same instance. Solutions from Gurobi(10s) are considered as the optimal benchmark in the simulations. The gap is computed by (26). $\aleph$ includes CoRaiS and other baselines. $\pi$ signifies the best solution obtained from the specific approach $b$ ($b\in\aleph$). $\hat{\pi}$ refers to the best solution generated by $Gurobi(10s)$. $L(\pi)$ refers to the predicted response time for all requests in the multi-edge cooperative computing system, and $L(\pi)$ is computed by (23). Gap-M is critical because the multi-edge cooperative computing system also require the real-time approach to produce a high-quality solution.
  • Time-M: average time taken to make scheduling decisions. Time-M is an important index to estimate whether the methods can provide effective decision in time.
  • Time-S: the standard variance of computing times to make scheduling decisions. Time-S can evaluate the stability of decision-making time.
TABLE V: Generalization Test Results on Lager-scale Instances
CoRaiS is trained on small-scale instances (10, 5, 100). Cost-M, Gap-M and Time-M: the lower the better.
(200, 100, 2k) (600, 300, 4k) (1000, 500, 6k) (2k, 1k, 10k) (2k, 1k, 15k)
Methods Cost-M Gap-M Time-M Cost-M Gap-M Time-M Cost-M Gap-M Time-M Cost-M Gap-M Time-M Cost-M Gap-M Time-M
Gurobi(5s) (0%) - - (0%) - - (0%) - - (0%) - - (0%) - -
Gurobi(10s) (3.1%) - - (0%) - - (0%) - - (0%) - - (0%) - -
CoRaiS(1k) 5.1795 1.0000 0.0600 5.8071 1.0000 0.3783 6.0991 1.0000 0.9563 7.1192 1.0000 3.1611 8.0548 1.0000 4.6019
Local 16.6937 3.2390 0.0403 14.5508 2.5340 0.1285 14.1081 2.3484 0.2129 13.7089 2.0241 0.6159 16.7721 2.2784 0.4401
Predicted 17.0419 3.3059 0.1293 14.5757 2.5398 0.3452 14.1772 2.3591 0.4828 13.7813 2.0354 1.2682 16.7845 2.2807 1.1119
Random(100) 13.0677 2.5316 0.0441 12.0275 2.0924 0.1426 11.8956 1.9807 0.2227 11.6695 1.7260 0.6598 14.4036 1.9613 0.6160
Random(500) 12.4988 2.4216 0.0573 11.6141 2.0199 0.2261 11.5509 1.9227 0.5283 11.3679 1.6806 1.7379 14.0588 1.9127 2.4922
Random(1k) 12.3202 2.3865 0.0796 11.4478 1.9912 0.4145 11.4235 1.9008 1.0049 11.2717 1.6671 3.2203 13.9004 1.8908 4.5946
CoRaiS(10) 5.2763 1.0186 0.0485 6.0588 1.0396 0.1547 6.4901 1.0604 0.2459 7.9634 1.1092 0.7364 9.2483 1.1436 0.6991
CoRaiS(100) 5.2053 1.0049 0.0597 5.9007 1.0139 0.1562 6.1954 1.0147 0.2724 7.4796 1.0441 0.7771 8.5213 1.0547 0.8522
CoRaiS(500) 5.1847 1.0009 0.0510 5.8259 1.0027 0.2175 6.1188 1.0035 0.5298 7.1911 1.0093 1.7753 8.1929 1.0183 2.6391
TABLE VI: Characteristics Validation Results
@@
CoRaiS is trained on scheduling problems (10, 5, 100).    CoRaiS(1k) is used to sample near-optimal solutions. LB: Load Balancing; WP: Workload Perception; HA: Heterogeneity Awareness; EReqN: the number of requests that are executed at each edges; LCost: the corresponding response time.
@@
A 21.5950 4.1647 19.7030 14.7180 11.9850 4.2231
B 19.6170 4.1958 20.2250 14.9676 14.8660 4.2165
C 19.5190 4.1825 20.0240 14.8704 21.5160 4.0770
D 19.6090 4.1935 20.0240 14.8709 25.8380 3.6357
E 19.6600 4.1953 20.0240 14.8709 25.7950 2.4347
VII-B Results Analysis
VII-B1 Analysis of Conventional Test Results